give solution
y=x-1
y=2x+3

Answer by mathbird(71)

y=x-1-----------eq1
y=2x+3----------eq2
from eq 1 ,the value of y is put in eq 2
x-1=2x+3
x-2x=3+1
-x=4
x=-4